OpenStudy (anonymous):

Need help solving the following for x, Sin(2x)-Tan(x)=0

OpenStudy (mathmale):

Suggestion 1: There is an identity for sin 2x which would be helpful here. Please look it up and use it. Suggestion 2: the tangent function can be re-written in terms of the sine and cosine functions.

OpenStudy (anonymous):

I know you can use the double-angle identity for sine. So it would be 2Sin(x)Cos(x)-SIn(x)/Cos(x). But were do i go from there.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

mult both sides by cos(x). but remember, \(x = \Large\frac{\pi}{2} \normalsize{+k\pi,\,\,k \in \mathbb{Z}}\) is not a solution (division by 0). then see what you get and if you can solve it.

OpenStudy (anonymous):

\[ \sin(2x) = 2\sin(x)\cos(x) \]\[ \tan(x) = \frac{\sin(x)}{\cos(x)} \]

OpenStudy (anonymous):

sin2x - tanx = 2sinxcosx - sinx/cosx = sinx(2cosx - 1/cosx) = 0 So sinx = 0, which happens when x = 0,180 or 2cosx = 1/cosx cosx = +/-(1/root2)

OpenStudy (anonymous):

x = 45, 135, 225, 315